OpenStudy (anonymous):

What is the exact value of tan 5pi/12?

OpenStudy (anonymous):

\[\frac{5\pi}{12}\] is half of \[\frac{5\pi}{6}\] i forget the half angle formula for tangent, but use that with \[\theta=\frac{5\pi}{12}\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

half angle formula is below http://home.windstream.net/okrebs/page103.html

OpenStudy (anonymous):

you can use \[\frac{\sin(\frac{5\pi}{6})}{1+\cos(\frac{5\pi}{6})}\]

OpenStudy (anonymous):

A. 2+ the square root of 3 B. 3.73 C.0.02 D.2- the square root of 3

OpenStudy (anonymous):

idk compute the number i wrote above

OpenStudy (anonymous):

it is alright but thanks anyways

OpenStudy (anonymous):

compute this number \[\frac{\sin(\frac{5\pi}{6})}{1+\cos(\frac{5\pi}{6})}\] that will give you the answer

OpenStudy (anonymous):

but how I don't know what you mean

OpenStudy (anonymous):

oh never mind it is A

OpenStudy (anonymous):

you need two numbers to put in that formula in the numerator you need \[\sin(\frac{5\pi}{6})\] and in the denominator you need \[\cos(\frac{5\pi}{6})\]